Model each problem with a quadratic equation. Then Solve.

The area of a triangular sail on a sailboat is 300ft squared. The ratioof the base of the sail to its height is 2:3. Find the dimensions of the sail.

Answer by khwang(438)

Let the height of the sail is x,
since base : = 2:3, so base = 2x/3,
Area = 1/2* base * height = 1/2* (2x/3) * x = x^2/3 = 300,
so x^2 = 900, x = 30 ft
and base = 2/3* 30 = 20 ft